# Stabilization of active tissue deformation by a dynamic morphogen gradient

### v1

@@ IV Localized source profile
In the following, we discuss how the linear stability changes when the source profile is not linear but instead localized to some region of finite width $\mathrm{{w}}$. This is motivated by what we know of developing biological tissues [1, 38, 33, 42]. In this section, we discuss a system of fixed size, while in section V, we discuss a freely deforming system.
As initial state, we use a source profile that is similar to a von Mises distribution, which is a generalization of a Gaussian to periodic boundary conditions:
@@
with $q_{0}=2\pi/L_{x}$. Here, $\mathrm{{w}}$ is a parameter that adjusts the width of the source region, and the second term in Eq. (20) ensures that $s_{0}(\bm{r})$ is zero at its minimum. As initial $c_{0}$ field, we choose the $c$ field that is stationary in the absence of flows for $s_{0}$ given by Eq. (20). In Fourier space:
@@
Thus, for $D=0$, morphogen and source profiles coincide, and examples for $s_{0}(x)=c_{0}(x)$ for different $\mathrm{{w}}$ values are shown in Figure 3A. Meanwhile, finite diffusion $D>0$ smoothens the morphogen profile $c_{0}(x)$ as compared to the source profile $s_{0}(x)$ (Figure 4A).
Figure 3: Stability of a fixed-size system with a localized source profile and a non-diffusive morphogen, $D=0$. (A) The localized source profile $s_{0}(x)$ from Eq. (20) for different values of the source width $\mathrm{{w}}$. (B,C) Growth rates $\omega_{m}$ of perturbation modes depending on their mean wavenumber $\langle n\rangle_{m}$ for a system with $\mathrm{w}=1/2\pi$, for (B) a gradient-contractile and a (C) gradient-extensile system. The eigenvalue problem was solved numerically as described in Appendix C.3. (D) Zoomed region of panel C displaying high-wavenumber perturbation modes with $\omega>0$, which we interpret as numerical artifacts, due to the spatial discretization (see main text, Appendix E.2, and panel G). (E-G) Morphohen field perturbation $\delta c(\bm{r})$ for three different eigen modes (as indicated in panels B and D). (E) $\delta c$ mode for small wavenumber $\langle n\rangle$, with positive growth rate in the gradient-contractile case. (F) Typical $\delta c$ mode with high $\langle n\rangle$ and high growth rate in the gradient-contractile case. (G) Typical $\delta c$ mode with numerically obtained positive growth rate in the gradient-extensile system. (H) Maximal growth rate $\omega_{\mathrm{max}}$ in gradient-contractile and -extensile systems depending on the numerical discretization, $N_{x}$. For gradient-extensile systems, this growth rate decreases as $\omega_{\mathrm{max}}\sim N_{x}^{-2}$. (I) Perturbation growth rate depending on the source width $\mathrm{{w}}$. Squares indicate $\omega_{\mathrm{max}}$, whereas triangles indicate the growth rate of perturbation modes in with small $\langle n\rangle$ (see main text and panel E).
@@
In this initial state, there are no flows in the system, $\bm{{\mathrm{v}}}_{0}=\bm{0}$. This is because the active stress in this case is parallel to the $x$ axis and homogeneous along the $y$ axis, because of incompressibility, and of the fixed system size, which together imply that there are no Stokes flows (Appendix B.2).
For a localized source, we study linear stability numerically. This is because the non-uniform gradient in the $c$ field creates non-constant coefficients in the linearized perturbation dynamics. As a consequence, the linearized dynamics in Fourier space now includes Fourier-space convolutions of the stationary profile $c_{0}$ and the morphogen perturbations $\delta c$ (Appendix C.2). Because of this, while the eigen modes of the linearized dynamics still grow or shrink exponentially with time, they are not single Fourier modes any more. We thus compute the perturbation eigen modes and their respective growth rates numerically.
For the case without diffusion, $D=0$, we plot in Figure 3B, C the growth rates $\omega_{m}$ of the different eigen modes $m$, where Figure 3B is for a gradient-contractile system and Figure 3C for a gradient-extensile system. For each eigen mode, the growth rates are plotted over their respective mean wavenumber $\langle n\rangle_{m}$, which we define as
$\langle n\rangle_{m}:=\frac{1}{2\pi}\,\frac{\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}\,q}}{\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}}},$ (22)
in dimensionless units (i.e. $L_{x}=L_{y}=1$), and where $\delta c_{m}(\bm{q})$ denotes the morphogen field part of eigen mode $m$ in Fourier space. For the gradient-contractile case, Figure 3B, we thus find that also for a non-uniform source the system is unstable, since there are many eigen modes with a positive growth rate $\omega_{m}>0$.
Surprisingly, Figure 3C seems to suggest that also gradient-extensile systems are unstable in this case, because there are modes with a positive growth rate (see magnified region in Figure 3D). However, these modes all have almost maximal wave number $\langle n\rangle_{m}\gg 1$ (Figure 3D, G). We thus wondered whether these positive growth rates might a numerical artifact, for instance due to the numerical discretization of the source and morphogen fields. To test this, we quantified the maximum of all growth rates $\omega_{m}$ for increasing numerical resolution in Figure 3H. We find that for the gradient-extensile case (green squares) the maximal growth rate decreases as $\omega_{\mathrm{max}}\sim N_{x}^{-2}$ with the numerical resolution in $x$ direction, $N_{x}$, while it remains largely constant for the gradient-contractile case (orange squares). This indicates that the positive growth rates of the large-$\langle n\rangle_{m}$ modes in the gradient-extensile case shown in Figure 3C inset are due to a numerical discretization artifact, without which the maximal growth rate in the gradient-extensile case would be zero. We independently confirm this further in appendix Appendix E.2.
Next, we study for $D=0$ how the system’s stability depends on the source width $\mathrm{{w}}$ (Figure 3I). The gradient-contractile system (orange symbols) is always unstable, where the squares in Figure 3I show the maximum growth rate, which decreases with source size, but approaches a positive value for $\mathrm{{w}}\rightarrow\infty$. As indicated by Figure 3B, this maximal growth rate can correspond to a mode $m$ with a large wavenumber $\langle n\rangle_{m}$ (Figure 3F). However, in biological tissues there is a natural upper wave vector cutoff, provided for instance by the cell size. Thus, to test in how far our results are still valid in the presence of some wave vector cutoff, we also indicate the growth rates of a small-wavevector mode. We selected the mode with $\langle n_{x}\rangle_{m}\neq 0$ and $\langle n_{y}\rangle_{m}\neq 0$ that has the smallest $\langle n\rangle_{m}$, where we used $\langle n_{i}\rangle_{m}:=\left(\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}\,q_{i}}\right)/\left(2\pi\sum_{\bm{q}}{\big{|}\delta c_{m}(\bm{q})\big{|}}\right)$. We find that also this small-wavevector mode has a positive growth rate for all source widths, which also approaches a finite value for $\mathrm{{w}}\rightarrow\infty$ (orange triangles in Figure 3E). Thus, we expect the material to be unstable even in the presence of some upper wave vecor cutoff. In contrast, the gradient-extensile system (green symbols) is always marginally stable, where the squares show again the maximum growth rate. The squares are open to indicate that we ignored any modes with large wavevecor and positive growth rates based on our discussion in the previous paragraph.
Figure 4: Effect of finite morphogen diffusion, $D>0$, on the stability of the fixed-size system with a localized source. Diffusion can not stabilize a gradient-contractile system. (A) Stationary morphogen profiles $c_{0}(x)$ for different diffusion coefficients, where $\mathrm{{w}}=1/2\pi$. (B) Perturbation growth rates of gradient-contractile systems for the three diffusion coefficients indicated in panel A. (C) Perturbation growth rates with increasing $D$. Squares indicate maximal growth rates $\omega_{\mathrm{max}}$ and triangles indicate growth rates of modes in the small $\langle n\rangle$ limit, $\omega_{\mathrm{s}}$.
Finally, we study how the system stability depends on the diffusion coefficient $D$. In this case, gradient-extensile systems are still marginally stable for any $D>0$ (Figure 4C). They are not further stabilized by diffusion for the same reason as for a linear source profile: any perturbation with $\phi=0$ in the source field neither induces any flows, nor can it “diffuse away”.
For gradient-contractile systems, diffusion decreases the growth rates of the large-wavevector modes (Figure 4B). Thus, for large enough diffusion constant, the system’s fastest growing mode becomes the small-wavevector mode. Hence, the maximum growth rate (orange squares in Figure 4C) overlaps with the growth rate of the small-wavevector mode (orange triangles) for $D\gtrsim 10^{-2}$. Remarkably, while for a linear source profile, a sufficiently high diffusion coefficient could bring down the maximal perturbation growth rate arbitrarily close to zero (Figure 2C, Eq. (19)), this is not possible any more for a localized source. Instead, for a localized source, the maximum perturbation growth rate is always larger than $\approx 2$ in units of the free deformation rate (Figure 4C).
Figure 5: We introduce co-deforming coordinates to study a deforming system, Eqs. (23) and (24). (A) In lab-frame coordinates, a localized source $s(\bm{r},t)$ is not stationary when the system is deforming affinely, because its width changes proportionally with the system dimension $L_{x}$ due to advection. (B) However, in co-deforming coordinates, the localized source $s(\bm{\mathrm{bar}{r}},t)$ remains stationary under affine deformation.
@@ V Freely deforming system
Here we study the linear stability of a system with a localized source $s_{0}$, Eq. (20), under free deformation. In this case, even for homogeneous shear deformation, the source profile $s_{0}$ is not stationary, but becomes distorted due to advection with the pure shear deformation (Figure 5A).
However, the source profile is stationary with respect to “co-deforming” coordinates. Following our earlier work [32], we define the co-deforming coordinates, $\bm{\mathrm{bar}{r}}=(\mathrm{bar}{x},\mathrm{bar}{y})$, based on the lab-frame coordinates $\bm{r}=(x,y)$ by rescaling both axes:
@@
where $l_{x}(t):=L_{x}(t)/L_{x}(0)$ is the box shear. Moreover, we also introduce a flow field in co-deforming coordinates, $\bm{\mathrm{bar}{{\mathrm{v}}}}(\bm{\mathrm{bar}{r}})$, which corresponds to non-affine flows that may occur in addition to the pure shear deformation of the periodic box (Appendix A.3). Thus, for $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$ the system deforms affinely. In this case, the source field $s_{0}(\bm{\mathrm{bar}{r}})$ is stationary with respect to the co-deforming coordinates (Figure 5B).
V.1 No morphogen diffusion, $D=0$
Without diffusion, the morphogen concentration given by $c_{0}(\mathrm{bar}{x})=s_{0}(\mathrm{bar}{x})$ is stationary with respect to the co-deforming coordinates, and the system deforms affinely, $\bm{\mathrm{bar}{{\mathrm{v}}}}_{0}=\bm{0}$. This is again due to the homogeneity of $c_{0}$ (and thus the active stresses) in $y$ direction combined with incompressibility (Appendix B.2).
Figure 6: Stability of a freely deforming system with a non-diffusive morphogen, $D=0$. (A) Time evolution of aspect ratios of a gradient-extensile system (green) and a gradient-contractile system (orange) starting from $L_{x}(0)/L_{y}(0)=0.04$ and $L_{x}(0)/L_{y}(0)=25$, respectively. (B) The corresponding absolute values of the shear rates of the systems in panel A. (C) Instantaneous perturbation growth rates over time for the systems in panel A. Square symbols display the maximal growth rates $\omega_{\mathrm{max}}$, whereas triangles display growth rates of small-wavenumber perturbations $\omega_{\mathrm{s}}$ (see Figure 3E). (D) Same as in panel C, where the perturbation growth rates are rescaled by the instantaneous shear rate.
System dimension dynamics. The dynamics of system dimension $L_{x}$, and thus box shear $l_{x}$, is given by Eq. (7). Thus, in the stationary state, the shear rate is essentially given by the integral over $(\partial_{x}c_{0})^{2}$. Because the morphogen profile, $c_{0}$, is stationary in co-deforming coordinates, we have that for varying system size, $(\partial_{x}c_{0})^{2}\sim 1/l_{x}^{2}$. So, using dimensionless units (subsection II.3), we can write
$\frac{1}{l_{x}}\,\frac{\mathrm{d}l_{x}}{\mathrm{d}t}=-\frac{\mathrm{sgn}(\alpha)}{l_{x}^{2}L_{x}^{2}(0)}.$ (25)
The solution of this is:
$l_{x}(t)=\sqrt{1-\frac{2\mathrm{sgn}(\alpha)t}{L_{x}(0)^{2}}}.$ (26)
Thus, gradient-extensile systems ($\alpha<0$) deform slower and slower due to the widening of the morphogen profile (green curves in Figure 6A,B). Meanwhile, gradient-contractile systems ($\alpha>0$) deform faster and faster over time due to the compression of the morphogen profile, until the deformation rate diverges at some critical time point $t_{\mathrm{crit}}=L_{x}^{2}(0)/2$ (orange curves in Figure 6A,B).
Perturbation growth rates. Without diffusion, the dynamics of source and morphogen perturbations, $\delta s$ and $\delta c$, are independent of the box shear perturbation, $\delta l_{x}$ (Eqs. (50) and (51) in Appendix C.1), but they still do depend on box shear $l_{x}(t)$ itself. Thus, to characterize the system’s stability over time, we can use our linear stability analysis for the fixed-size system (section IV and Figure 3), were we use the time-dependent $l_{x}$ according to Eq. (26). The results are shown in Figure 6C: As in section IV and Figure 3, the gradient-contractile systems are always unstable, where even their smallest-wavevector perturbation grows (orange symbols), while the gradient-extensile system is always marginally stable with a decaying smallest-wavevector perturbation (green triangles). We moreover observe that for gradient-contractile systems, the perturbation growth rates become substantially larger at late times, where $l_{x}$ is small. Meanwhile, for gradient-extensile systems, the magnitude of the decay rates is substantially larger at earlier times, where again $l_{x}$ is small. This is likely due to the stronger localization of the source region due to the smaller $l_{x}$; indeed, in Figure 3I we showed that the magnitudes of the perturbation growth rates are larger for smaller source sizes $\mathrm{{w}}$. To better compare to the box shear rate, in Figure 6D we normalize these perturbation growth rates by the instantaneous box shear rate from Figure 6B. We find that the perturbation growth rate magnitudes normalized by the instantaneous shear rates vary much less over time, and is typically on the order of $10$, both for gradient-contractile and -extensile systems. They only become a little smaller for systems with small $l_{x}$ (i.e. late times for gradient-contractile systems, and early times for gradient-extensile systems). Figure 6D allows to predict the expected perturbation growth rates for a biological system of known instantaneous free deformation rate.
V.2 Finite morphogen diffusion, $D>0$
@@
In a fixed-size system with the source field $s_{0}$ given by Eq. (20) and with $D>0$, the interaction between production, diffusion, and degradation allowed for existence of a stationary state (section IV). However, when such a system is allowed to freely deform, it does not reach a stationary state – not even in co-deforming coordinates. This is because diffusion as observed in co-deforming coordinates depends on box shear $l_{x}$ through the transformation rules in Eqs. (23) and (24). Thus, as a consequence of the permanently changing $l_{x}$, there is no stationary state in the freely deforming case.
@@
Figure 7: Stability of a freely deforming system with a diffusive morphogen, where we set $D=10^{-2}$ and in the limit $\mathrm{{w}}\rightarrow\infty$, where the source has a cosine shape. (A) Aspect ratio over time for the gradient-extensile (green) and contractile (orange) systems, starting at initial aspect ratios of $L_{x}(0)/L_{y}(0)=1/225$ and $L_{x}(0)/L_{y}(0)=25$, respectively. (B) The corresponding absolute values of the shear rates of the systems in panel A. Inset: zoomed around the sharp drop in the gradient-contractile case. (C) Illustration of the affine diffusion-shear instability: diffusion reduces the amplitude $\hat{c}$ of the morphogen field, but a larger $L_{x}$ decreases the effectiveness of diffusion. Thus, a larger $L_{x}$ leads to a larger $\hat{c}$, which in turn increases active shear and thus $L_{x}$. (D) Local Lyapunov exponents (LLE) for the gradient-extensile system over time. Green crosses indicate the numerically computed LLE of the affine diffusion-shear perturbation with the largest growth rate, and the blue curve displays the analytical prediction according to Eq. (88) in Appendix F.2.1. Green squares denote the maximal LLE of all other perturbations. (E) The same results as in panel D, but for the gradient-contractile system.
System size dynamics. In a gradient-extensile system, we observe that the aspect ratio increases over time as expected (green curve in Figure 7A). However, the shear rate first increases and then decreases (green curve in Figure 7B), in contrast to the case without diffusion, where it monotonically decreases (Figure 6B). For gradient-contractile systems, we observe a much more erratic system size behavior (orange curves in Figure 7A,B). In particular, the shear rate shows a transient, but pronounced peak before quickly switching to much slower dynamics (Figure 7B inset). We discuss the reason for these $l_{x}$ dynamics in detail using scaling arguments in Appendix F.1.
@@
We first briefly discuss the local Lyapunov exponents in the limit of no diffusion, $D=0$ (previous section, subsection V.1). For $D=0$, the linearized dynamics of source and morphogen, Eqs. (50) and (51) (Appendix C.1), do not depend on the perturbation of the box shear, $\delta l_{x}$. Thus, the perturbation dynamics of $s$ and $c$ decouples from that of $l_{x}$, and most of the Lyapunov spectrum is what we already discussed in the previous section. The remaining eigenvalue of the local Lyapunov spectrum is given by the system size perturbation dynamics:
@@
The prefactor in this equation is positive for a gradient-contractile system and negative for a gradient-extensile system. Hence, for $D=0$, the gradient-contractile system has positive local Lyapunov exponents, i.e. the system dynamics strongly depends on the initial conditions. Meanwhile for gradient-extensile systems, the maximal local Lyapunov exponent is zero, and so the system dynamics depends less strongly on the initial conditions.
@@
For the gradient-contractile case, we found for $D=0$ several modes with a positive growth rate. For $D>0$ this is still true (orange squares in Figure 7E), and – like in the gradient-extensile case – we again have an eigen mode of the Jacobian that is an affine diffusion-shear perturbation, i.e. it does not affect source field and non-affine flows (black crosses in Figure 7E compare predicted growth rate marked by the blue line).
Finally, we studied how the instability due to the affine diffusion-shear perturbation is changed for the general case of finite source width $\mathrm{{w}}$, and we find that this perturbation has a positive growth rate whenever the order of magnitude of the diffusion length becomes larger than the source width, $\lambda\gtrsim\mathrm{{w}}$ (Appendix F.2.2).

### v2

@@ IV Localized source profile
We next discuss how the linear stability changes when the source profile is not linear but is instead localized to some region of finite width $\mathrm{{w}}$. This is motivated by what we know about developing biological tissues [1, 41, 38, 45]. In this section, we discuss a fixed-size system $L_{x}=L_{y}=1$, while in the next one (section V) we discuss a freely deforming system.
[…]
As initial state, we use a source profile defined by a von Mises distribution, which is a generalization of a Gaussian to periodic boundary conditions:
@@
with $q_{0}:=2\pi/L_{x}$. Here, $\mathrm{{w}}$ is a parameter that adjusts the width of the source region, the second term in Eq. (21) ensures that $s_{0}(\bm{r})$ is zero at its minimum, and $s_{b}$ is a prefactor whose value is fixed such that the box deformation rate is normalized to one (Eq. (75) in Appendix E.1). As initial signaling field, we choose the $c$ field that is stationary in the absence of flows for the source given by Eq. (21). In Fourier space:
@@
Thus, for $D=0$, signaling and source profiles coincide, and examples for $s_{0}(x)=c_{0}(x)$ for different $\mathrm{{w}}$ values are shown in Figure 3A. Meanwhile, finite diffusion $D>0$ smoothens the signaling profile $c_{0}(x)$ as compared to the source profile $s_{0}(x)$ (Figure 3B).
@@
In this initial state, there are no flows in the system, $\bm{{\mathrm{v}}}_{0}=\bm{0}$. This is because the active stress in this case is parallel to the $x$ axis and homogeneous along the $y$ axis, because of incompressibility, and because of the fixed system size, which together imply that there are no Stokes flows (Appendix B.2).
Numerical simulations of the full dynamics, Eqs. (10)–(13), are shown in Figure 3C,D and in Movies S1, S2 (shown is the source field). We find that with chosen parameters ($\mathrm{w}=1/2\pi$, $k_{\mathrm{d}}=1$, and $D=10^{-2}$), the gradient-extensile system appears to be stable until simulation time $t=3$ (Figure 3C), yet an instability becomes apparent at much later times (Figure 5A in Appendix E and Movie S1). Meanwhile, the gradient-contractile system shown an instability already early on (Figure 3D and Movie S2).
For a more systematic characterization of the system’s stability, we perform again a linear stability analysis. While the eigenmodes of the linearized dynamics still grow or shrink exponentially with time, they can only be represented by a superposition of several spatial Fourier modes (Appendix C.2). We thus compute the perturbation eigenmodes and their respective growth rates numerically (Appendix C.3). Note that due to the numerical Fourier-space convolutions, we also find aliasing effect arising from a finite spatial discretization, which we treat using standard methods in the field (Appendix C.3).
We find again that gradient-contractile systems are always linearly unstable (Figure 3E,F,H,J). Specifically, systems are more unstable when the source width is below $\mathrm{{w}}\lesssim 0.06L_{x}$ (Figure 3F,H). Moreover, increasing the diffusion coefficient $D$ decreases the perturbation growth rate, but it never reaches zero. Instead, in contrast to the linear gradient (section III), the maximum growth rate converges to a finite constant in the limit of large diffusion, $D\rightarrow\infty$ (Figure 6, Appendix E.2). It converges to a finite constant, because we non-dimensionalize rates with respect to the free deformation rate, which also decreases as diffusion $D$ becomes larger due to an increased smoothing of $c$ (Appendix E.2). This is different from the linear-gradient scenario (section III), where a constant signaling gradient, and thus a constant free deformation rate is imposed by the boundary conditions.
For gradient-extensile systems, the situation is more complex than for a linear gradient. Specifically, we find that the stability depends on the aspect ratio of the system. For aspect ratios $L_{x}/L_{y}$ below $2\dots 3$, the system can be unstable, while it is numerically marginally stable for aspect ratios above (Figure 3E,G). Here, by “numerically marginally stable” we mean that we numerically cannot distinguish between marginally stable and numerically small, but finite growth rates – indeed, this distinction is also irrelevant in practice for a real biological system. For aspect ratios, between $\approx 1$ and $2\dots 3$, the system is effectively unstable at intermediate source widths $\mathrm{w}/L_{x}$ between $0.04$ and $0.2$ and otherwise numerically marginally stable (white regions in Figure 3G). For even lower aspect ratios, we only observe numerical marginal stability for very narrow sources with $\mathrm{w}/L_{x}\lesssim 0.04$ (Figure 3F,G). The type of perturbation creating the most unstable mode appears to be always qualitatively the same: the local width of source and signaling is modulated by the flows created by the signaling perturbation (Figure 5B). Finally, for given aspect ratio and source width, the system is generally more stable for small signaling length scales $\sqrt{D/k_{\mathrm{d}}}$ (Figure 3I). In the large-diffusion limit, the maximum perturbation growth rate converges to a constant for the same reason as in the gradient-contractile case (Figure 6, Appendix E.2).
Note that even if gradient-extensile systems are formally unstable in a substantial part of the parameter space, for aspect ratios close to one the maximal perturbation growth rates can be several orders of magnitude smaller as compared to the gradient-contractile case. For instance, the maximal growth rate for the example simulation in Figure 3C, Figure 5A, and Movie S1 is positive, but very small, $\omega_{\mathrm{max}}=1.9\times 10^{-3}$ (as compared to $\omega_{\mathrm{max}}=4.2$ in the gradient-contractile example in Figure 3D).
Taken together, the system is stable when it is more elongated in the direction of the gradient, or when the signaling length scale $\sqrt{D/k_{\mathrm{d}}}$ is small.
@@ V Freely deforming system
Here we study the linear stability of a system with a localized source $s_{0}$, Eq. (21), under free deformation of the periodic box.
Even for homogeneous shear deformation, the source profile $s_{0}$ is not stationary, but becomes distorted due to advection with the box deformation (Figure 4A). To address this, following our earlier work [35], we define the co-deforming coordinates, $\bm{\mathrm{bar}{r}}=(\mathrm{bar}{x},\mathrm{bar}{y})$, based on the lab-frame coordinates $\bm{r}=(x,y)$ by rescaling both axes:
@@
where $l_{x}(t):=L_{x}(t)/L_{x}(0)$ is the box shear. Moreover, we also introduce a flow field in co-deforming coordinates, $\bm{\mathrm{bar}{{\mathrm{v}}}}(\bm{\mathrm{bar}{r}})$, which corresponds to non-affine flows that may occur in addition to the pure shear deformation of the periodic box (Appendix A.3). Thus, for $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$ the system deforms affinely, and the source field $s_{0}(\bm{\mathrm{bar}{r}})$ is stationary with respect to the co-deforming coordinates (Figure 4B).
@@
Yet, in the general case of a finite diffusion constant, the signaling field $c$ does not reach a stationary state – not even in co-deforming coordinates. This is because diffusion as observed in co-deforming coordinates depends on box shear $l_{x}$ through the transformation rules in Eqs. (23) and (24). Thus, as a consequence of the permanently changing $l_{x}$, there is no stationary state in the freely deforming case. Hence, no standard linear stability analysis is possible – we instead analyze the local Lyapunov exponents of the system.
Figure 4: Stability of a freely deforming system with a diffusive signaling molecule, studied in co-deforming coordinates. (A) In lab-frame coordinates, a localized source $s(\bm{r},t)$ is not stationary when the system deforms affinely, since its width changes proportionally with the system dimension $L_{x}$ due to advection. (B) However, in co-deforming coordinates, Eqs. (23) and (24), the localized source $s(\bm{\mathrm{bar}{r}},t)$ remains stationary under affine deformation. (C) Illustration of the affine diffusion-shear instability: diffusion reduces the amplitude $\hat{c}$ of the signaling field, but a larger $L_{x}$ decreases the effective diffusion coefficient observed in co-deforming coordinates. Thus, a larger $L_{x}$ leads to a larger $\hat{c}$, which in turn increases active shear and thus $L_{x}$. (D) Local Lyapunov exponents (LLE) for the gradient-extensile system over time. Green crosses indicate the numerically computed LLE of the affine diffusion-shear perturbation with the largest growth rate, and the blue curve displays the analytical prediction according to Eq. (85) in Appendix F.2.1. Green squares denote the maximal LLE of all other perturbations. (E) The same results as in panel D, but for the gradient-contractile system.
@@
In the limit of no diffusion, $D=0$ (previous section, section IV), the linearized dynamics of source and signaling, Eqs. (48) and (49) (Appendix C.1), do not depend on the perturbation of the box shear, $\delta l_{x}$. Thus, the perturbation dynamics of $s$ and $c$ decouples from that of $l_{x}$, and most of the Lyapunov spectrum corresponds to the eigenvalue spectrum of the fixed system (section IV). The remaining eigenvalue of the local Lyapunov spectrum is given by the system size perturbation dynamics:
@@
The prefactor in this equation is positive for gradient-contractile systems and negative for gradient-extensile systems. Hence, for $D=0$, the gradient-contractile system has positive local Lyapunov exponents, i.e. the system dynamics strongly depends on the initial conditions. Meanwhile, for gradient-extensive systems, the maximal local Lyapunov exponent is given by the maximal eigenvalue for the fixed-size system (section IV).
@@
We first discuss the gradient-extensile case, where we find here that the maximal perturbation growth rate can be positive (blue crosses in Figure 4D). However, at later times, when $l_{x}$ becomes larger (green curves in Figure 7B, subsection F.1), the maximal growth rate becomes zero (green squares in Figure 4D). This fastest growing perturbation mode (blue crosses in Figure 4D), which we call here “affine diffusion-shear perturbation”, is different from the active scalar matter instability [53, 35]. Specifically, it does not affect the non-affine part of the flow field, $\bm{\mathrm{bar}{{\mathrm{v}}}}$ or the source field $s$. Moreover, the signaling profile keeps its cosine shape, and only its amplitude $\hat{c}$ and the box shear $l_{x}$ are affected by the perturbation. The positive growth rate of this mode arises from the fact that diffusion diminishes the amplitude of the signaling cosine profile, $\hat{c}$. Thus, increasing $l_{x}$ by some positive amount $\delta l_{x}$ decreases the effect of diffusion, which acts to increase $\hat{c}$ and thus the integrated squared signaling gradient, which in turn acts to further increase $l_{x}$ (Figure 4C). In Appendix F.2.1, Eq. (85), we obtain an analytical prediction for its growth rate (light blue curve in Figure 4D), which matches our numerical results (blue crosses). From our analytical prediction, we find that this instability appears as long as the diffusion length $\lambda:=\sqrt{D/k_{\mathrm{d}}}$ is larger than $L_{x}$, or more precisely $\lambda>L_{x}/2\pi\sqrt{3}$ (Appendix F.2.1). Remarkably, for large diffusion, this instability does not disappear even when introducing an arbitrary non-linearity into the active stress (Appendix G).
For the gradient-contractile case, we find several modes with a positive growth rate (orange squares in Figure 4E). Moreover, we again have an eigenmode of the Jacobian that is an affine diffusion-shear perturbation (crosses in Figure 4E, compare to the predicted growth rate marked by the light blue curve).
Finally, in Appendix F.2.2 we study the affine diffusion-shear instability for the general case of a finite source width $\mathrm{{w}}$. We find a positive growth rate whenever the order of magnitude of the diffusion length becomes larger than the source width, $\lambda\gtrsim\mathrm{{w}}$ (Appendix F.2.2).